Determine the final temperature of the liquid when you mix:

m1 = 1.00 kg water at t1 = 10.0 °C

and

m2 = 2.00 kg water at t2 = 40.0 °C.

Assume that no heat is lost to the environment.

A numerical value is required for credit.

SHOW YOUR WORK.
Answer #1

let the final temperature be t

heat lost by hot water = heat gained by old water

mc*dt for hot water = mc*dT for cold water

1*c*(40-t) = 2*c*(t - 10)

40 - t = 2t - 20

3t = 60

t = 20 oC

so the final temperature is 20o C